Bats use the Doppler effect to determine their location. A bat flies toward a wall at a speed of 15 m/s while emitting sound of frequency 35 kHz. What is the frequency received by the bat? (hint: 2 doppler effects occur!)
Answer #1

let the speed of sound in air is 300 m/s

the relative speed between them is 300 - 15 = 285 m/s


wavelength lambda = v /f

lambda = 285/ 35 X 10^3


but to a stationary observer at the wall

this incoming sound would have a frequency of

300 X 35 X 10^3/(300-15) = 36.84 kHz

the waves bounce off the wall at a frequency of 36.84 and head back to the bat

now the speed is 300 + 15 = 315 m/s

frequency f = 315 X 36.84 / 300 = 38.68 kHz

So the beat frequency heard by the bat on inbound

38.68 kHz - 35 kHz = 3.68 kHz


answered by: Sublower